The domain of the function define by f(x) = √[Inx/ (Inx-1)] is?



1
Expert's answer
2021-10-26T17:54:38-0400

The first restriction:


"x>0"

since it is not possible to have "x\le0" under the "\ln".

The second restriction:


"\ln x - 1\ne0\Rightarrow x\ne e"

since the denominator can not be zero.

The third restriction:


"\dfrac{\ln x}{\ln x - 1}\ge0"

since there can not be a negative quantity under the square root sign. The last restriction is satisfied if:


"\ln x\ge0\space\text{and}\space \ln x-1>0\\
\ln x\ge0\space\text{and}\space\ln x>1\\
x\ge1 \space\text{and}\space x>e\\
x>e"

or


"\ln x\le0\space\text{and}\space \ln x-1<0\\
\ln x\le0\space\text{and}\space\ln x<1\\
0<x\le1 \space\text{and}\space 0<x<e\\
0<x\le1"

Finaly, for the third restriction have:


"x\in (0;1]\cup(e;+\infty)"

Combining these three restrictions together, find the domain:


"x\in (0;1]\cup(e;+\infty)"


It is so, since the third restriction already contains the first and the second ones.


Answer. "x\in (0;1]\cup(e;+\infty)".